# Accreditation Against Limited Adversarial Noise

### v2

@@ Abstract
I present an accreditation (a variety of quantum verification) protocol where error is assumed to be adversarial (in contrast to the IID assumption used in previous accreditation protocols) - albeit slightly modified to more closely resemble physical reality - with no diminution in efficiency or suitability for near-term usage.
@@ I Introduction
Accreditation provides an efficient and scalable method for quantifying the the quality of quantum computations, without trusting any aspect of the computations. This is a vital requirement for using quantum computers in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment - known as noise [2] - that induce erroneous operators in a computation; so some measure of whether the outputs of quantum computations are usable is needed.
I note that accreditation is not the only method of evaluating quantum computations or quantum computers. One such class of protocols is randomized benchmarking [3, 4, 5, 6]. These protocols do not aim to measure the how well a quantum computer can implement a specific computation - e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7] - but rather how well it performs benchmarking computations - which usually comprise of a series of gates that, if implemented with no error, return a known measurement outcome with certainty - and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented and if data obtained from quantum devices can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and - via some machinations - decide whether to accept or reject the returned outputs. This is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as in Def. 1) which accreditation protocols provide.
@@
For any circuit, $\mathcal{C}$, and an execution of it, $\tilde{\mathcal{C}}$, the ideal-actual variation distance of that execution (denoted as $\nu\big{[}\tilde{\mathcal{C}}\big{]}$) is the variation distance between the probability distribution the execution would sample from if there were no error (i.e. the ideal case) - which I denote by $\mathcal{O}_{I}\big{[}\mathcal{C}\big{]}$ and the distribution it actually samples from.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework of delegated computation [12] (i.e. is based on Alice and Bob interacting and precluding error typically [^1] There exists verification protocols that do not assume there is no error in some particular aspect of a computation, using two quantum devices [13, 14, 15], but here I focus on single-device protocols. either in state preparation [16, 17, 18] or measurement [19, 20, 14, 21]); accreditation has instead tended to base its assumptions on the physics within quantum computers. By turning from the cryptographic assumptions of verification protocols [8], from which it originates, to physics-based assumptions; accreditation has been able to develop assumptions, and then protocols [22] - using those assumptions - to initiate a vein of research that has, thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [23], and led to the first method for quantifying the quality of the outputs of quantum analogue simulation [24].
@@
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^2] Meaning in each execution of a circuit, the error is represented by identical and independent - across multiple executions - completely positive trace preserving maps acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [23] to use/assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [12], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in - still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge and abilities. This permits me to improve upon the IID CPTP error model, in Ref. [23], while retaining validity in all situations where it was valid. The result is a protocol with less stringent assumptions, loosening the requirements on the physical devices used to implement them.
@@
Table 1: Table showcasing the upgrade presented in this paper as a comparison between the assumptions in Ref. [23] and the protocol herein. Note that GI is shorthand for Gate-Independent (see Sec. II.4).
This is achieved through exploiting limits placed on Bob (based on experimental realities - that the probability of error varies little in executions of the same circuit, on the same hardware, in the same environment - and inspired by single-qubit gates experiencing gate-independent (GI) error in pre-existing accreditation protocols [22, 23]) allowing methods to overcome the factors that prohibited adversarial noise in extant accreditation protocols.
This paper proceeds, from here, with an introduction to the problem setting - in Sec. II.3 - and its justification - in Sec. II.4. I then present my main result: the adaptation of accreditation protocols to the newly-developed problem setting / model of error in Sec. III. The paper then concludes - in Sec. IV - with a discussion of further possible developments.
@@ II Problem Setting of This Paper
As this manuscript bridges the gap between the physics-based assumption model of previous accreditation papers and the cryptographic setting that is more typical of verification protocols, I take an approach that I hope will satisfy both tribes: below I present the problem setting - where Alice, Bob, and their respective objectives are introduced - the crux of which is an protocol (Protocol 1) for how circuits are executed (which Alice and Bob both participate in), and then the physics of errors occurring in a circuit execution are used to support and justify the problem setting, for those who prefer a physics-based error model.
The foundation of my problem setting [^3] Which can be seen as akin to the error model in previous accreditation protocols. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [23]. This is justified as any map from and to density matrices is a CPTP map. Hence, any erroneous operation must be a CPTP map and so any erroneous map may be written as the errorless/ideal operation either followed or preceded by a CPTP map - as unitaries are CPTP maps.
II.1 Introduction to Alice and Bob
With its foundations established, I now present my problem setting, which is an adaption of the cryptographic setting [22]. In the cryptographic setting, Alice is attempting to get Bob to perform a sampBQP [25, 26, 27] quantum computation [^4] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). and provide her with the true results, while only having the ability to perform polynomially-bounded classical computation herself. Bob is computationally unbounded so, if he wished, could implement the computation Alice requires but his aim is instead to trick Alice by providing her with the results of an incorrect computation and to have her believe he has acted honestly i.e. that the results she receives are the outputs of the computation she wanted to be performed when they are not.
II.2 Problem Setting: Preliminaries and Definitions
My specific problem setting is slightly modified from the traditional cryptographic setting, mainly through limitations on Bob. These limitations are entirely contained in how the circuits Alice requests be executed - and Bob tries to corrupt - are executed, which is presented in Protocol 1 in Sec. II.3, but Protocol 1 first requires the concept of redaction (introduced in Def. 2) and a related concept (in Def. 3).
@@ II.2 Problem Setting: Preliminaries and Definitions
A gate in a circuit is said to be redacted if the gate’s position in the circuit - i.e. when it is applied and to which qubits - is specified but the operator it represents is not. For an example of a circuit with a redacted gate, see Fig. 1.
Figure 1: An example circuit with a redacted gate (the block in the middle of the circuit). Note how the location of the gate - in terms of time and which qubits are affected - is depicted but which operation the gate represents is hidden.
@@
If given a circuit with all single-qubit gates redacted, the set of all circuits - with no redactions - that the given circuit (with redactions) could be (if the redactions were removed) is called its redaction class. For an example of circuits in the same redaction class, see Fig. 2.
@@
Figure 2: Example circuits within the same redaction class, where the redaction class corresponds to the redacted circuit in Fig. 1. Note that the circuit in Fig. 1 could be any of these if the redaction were removed.
@@
The second and final round of terminology required for the presentation of Protocol 1 is related to how I specify the CPTP maps that are modelling the error in a single execution of a circuit. This is performed via CPTP lists and they are defined in Def. 4.
@@
However, a CPTP list is just a list of CPTP maps. In order to to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps. In this case the CPTP list is said to fit the circuit.
@@
A CPTP list fits a circuit if it may be used to describe error  [^5] Any resulting error - including no error - is valid. It need not be in any way physically justified or correspond to a particular quantum computer’s typical error. in an execution of that circuit.
A CPTP list achieves this by containing exactly one CPTP map corresponding to each location in the circuit where error may occur (These locations are: immediately after gates, immediately after state prep, and immediately before measurement)  [^6] The mapping between the CPTP maps in the CPTP list and the locations error occurs in the circuit is made using the ordering of the CPTP list and matching this with an ordering applied to the locations in the circuit where error may occur..
@@
For an example of a CPTP list determining the error in a circuit execution, see Fig. 3.
@@
Note that notation is slightly abused to display CPTP maps as gates (depicting error) in a circuit, to remedy this slightly they are highlighted in red to denote they are not unitaries but are CPTP maps.
@@
An important feature to note about CPTP lists and the circuits it fits is mentioned - without proof - in Lemma 1. Proof is omitted as Lemma 1 follows from the above discussion.
@@
The error in an execution of a circuit is entirely and exclusively determined by the CPTP list used to describe its error (and the outcomes of any stochastic processes that may occur in the application of the error described by the CPTP list  [^7] As a CPTP map may describe the stochastic application of operations.).
@@
The probability of error in the execution of a circuit, due to a CPTP list, is the probability that the execution does not provide a sample from the same distribution that executing the circuit as intended (i.e. with no error applied) would. Due to Lemma 1, the probability of error is entirely determined by the CPTP list used to represent the error in the execution. For example, using $\Xi_{1}=[\xi_{1},\xi_{2},\xi_{3}]$ from Fig. 3, if, for all density matrices, $\rho$:
$\xi_{1}(\rho)$ $=\dfrac{1}{2}\rho+\dfrac{1}{2}\hat{X}\rho\hat{X}$ (1)
$\xi_{2}(\rho)$ $=\dfrac{1-0.01}{2}\rho+\dfrac{1+0.01}{2}\hat{X}\rho\hat{X}$ (2)
$\xi_{3}(\rho)$ $=\dfrac{1+0.01}{2}\rho+\dfrac{1-0.01}{2}\hat{X}\rho\hat{X},$ (3)
then the probability of error in the execution of a circuit where these CPTP maps (i.e. $\Xi_{1}$) determine the error is the probability that any of $\xi_{1}$, $\xi_{2}$, or $\xi_{3}$ cause error. In this case, the probability of error is $\approx 0.875$. This quantity is not always meaningful if the error in an execution is not stochastic Pauli error, therefore, for all the traps and target circuits used throughout this paper, the probability of error is defined to be the probability error occurs once the error has been effectively reduced to - and can be considered as - stochastic Pauli error (by the Pauli twirling, defined in Def. 8, and present in all traps and targets considered herein).
@@
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (henceforth denoted as a PSCLβ) is a CPTP list such that for any CPTP list in the set, the probability of error in any circuit execution it determines the error for is within the interval:
@@
for some $P_{0}\in[0,1]$. Additionally, all CPTP maps in the PSCLβ must be of the same size, so if one fits a given circuit all CPTP maps in the PSCLβ do.
For example consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 6). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as a PSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this PSCLβ is $0.5$).
With both PSCLβ and redaction classes defined, they may be related via the below Lemma 2. As will be seen in Protocol 1, Alice chooses a redaction class (implicitly, by choosing a subset of one to execute) and Bob chooses a PSCLβ to determine the error in those executions. Hence, PSCLβ and redaction classes are - in some sense - duals of each other. This duality is further expounded apon by Lemma 2.

### v3

@@ Abstract
I present an accreditation protocol (a variety of quantum verification) where error is assumed to be adversarial (in contrast to the assumption error is implemented by identical CPTP maps used in previous accreditation protocols) – albeit slightly modified to reflect physically motivated error assumptions. This is achieved by upgrading a pre-existing accreditation protocol (from [S. Ferracin et al. Phys. Rev. A 104, 042603 (2021)]) to function correctly in the face of adversarial error, with no diminution in efficiency or suitability for near-term usage.
@@ I Introduction
Accreditation is a type of verification that provides an efficient and scalable method for quantifying the quality of specific quantum computations, without trusting any aspect of those computations. Verification is a vital requirement for using quantum computers – and demonstrating quantum advantage – in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment (known as noise [2]) that induce erroneous operators (known as error) in a computation.
I note that accreditation is not the only method of evaluating quantum computations or quantum computers. One class of alternative protocols is randomized benchmarking [3, 4, 5, 6]. Randomized benchmarking protocols do not aim to measure how well a quantum computer can implement a specific computation (e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7]) but rather how well it performs benchmarking computations (which usually comprises a series of gates that, if implemented with no error, return a known measurement outcome with certainty) and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented and if specific data obtained from quantum computers can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and – via some machinations – decide whether to accept or reject [^1] On the grounds that noise has ruined the computation. the returned outputs. This binary measure is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as in Def. 1) which accreditation protocols provide.
@@
For any circuit, $\mathcal{C}$, and an execution of it, $\tilde{\mathcal{C}}$, the ideal-actual variation distance of that execution (denoted as $\nu\big{[}\tilde{\mathcal{C}}\big{]}$) is the variation distance between the probability distribution the execution of the circuit would sample from if there were no error (i.e. the ideal case) and the probability distribution it ($\tilde{\mathcal{C}}$) actually samples from.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework/setting of delegated computation [12] (i.e. is based on two characters Alice and Bob, each with their own opposing aims) interacting and precluding error typically [^2] There exist verification protocols that do not assume there is no error in some particular aspect of a computation, using two non-communicating quantum computers [13, 14, 15], but here I focus on single-device protocols. either in state preparation [16, 17, 18] or measurement [19, 20, 14, 21]); accreditation has instead tended to base its assumptions on the experimentally-observed physics within quantum computers. By turning from the cryptographic and trust-based assumptions of verification protocols [8] – from which it originates – to physics-based assumptions, accreditation has been able to develop assumptions and then protocols [22] – using those assumptions – to initiate a vein of research that has thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [23], and led to the first methods for quantifying the quality of the outputs of quantum analogue simulations [24, 25].
@@
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^3] Meaning in each execution of a circuit, the error is represented by identical and independent – across multiple executions – completely positive trace preserving maps (as defined in Def. 5) acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [23] to assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [12], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in – still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge (of the circuits to be executed) and abilities (to influence the execution of circuits) by adding a new, impartial, character, Robert, who actually perfroms all quantum computations. This upgrade is summarized in Table 1 and improves upon the IID CPTP error model, in Ref. [23], while retaining validity in all situations where it was valid before. The result is a protocol with less stringent assumptions, loosening the requirements on the physical computers used to implement it.
@@
Table 1: Table showcasing the upgrade presented in this paper as a comparison between the assumptions in Ref. [23] and herein. Note that GI is shorthand for Gate-Independent (see Sec. II.4).
This is achieved through exploiting limits placed on Bob, and enforced by Robert, (based on experimental realities – that the probability of error varies little between executions of the same circuit, on the same hardware, in the same environment – and inspired by single-qubit gates experiencing gate-independent (GI) error in pre-existing accreditation protocols [22, 23]) allowing for methods to overcome the factors that prohibited an upgrade to assuming adversarial noise in previously extant accreditation protocols.
This paper proceeds, from here, with an introduction to the adversarial problem setting used throughout this paper – in Sec. II.3 – and its justification – in Sec. II.4. I then present my main result: the adaptation of accreditation protocols to the newly-developed problem setting in Sec. III. The paper then concludes – in Sec. IV – with a discussion of further possible developments.
@@ II Problem Setting of This Paper
As this manuscript bridges the gap between the physics-based assumption model of previous accreditation papers and the cryptographic setting that is more typical of verification protocols, I take an approach that I hope will satisfy both tribes: below I present the problem setting (summarized in Table 2) – where Alice, Bob, Robert and their respective objectives are introduced – the crux of which is a protocol (Protocol 1) for how circuits are executed (which Alice and Bob both participate in but neither actually implement the computation themselves; instead, Robert – who is fanatically and exclusively devoted to correctly performing his role in Protocol 1 – does), and then the physics of errors occurring in a circuit execution are used to support and justify the problem setting, for those who prefer a physics-based error model.
The foundation of my problem setting [^4] Which can be seen as akin to the error model in previous accreditation protocols [22, 23, 24, 25]. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [23] and depicted in Fig. 3. CPTP maps are as defined in Def. 5, which first requires several definitions.
The first of these definitions is Def. 2.
@@
Any linear operator, $\hat{A}$, acting on any Hilbert space is positive if, for any element, $|x\rangle$, in the Hilbert space:
$\langle x|\hat{A}|x\rangle\geq 0.$ (1)
Def. 2 then enables me to define a positive map, in Def. 3.
@@
A map is positive if positive operators are mapped exclusively to positive operators by it. If, $\forall N\in\mathbb{N}$, $\Phi\otimes I_{N}$ is positive (where $I_{N}$ is the identity on $N$ qubits), then $\Phi$ is completely positive.
With one half of the attributes of CPTP maps defined, I turn to the other half, and define trace-preserving maps in Def. 4.
Definition 4.
A map, $\Phi$, is trace-preserving if for any density matrix, $\rho$:
$\operatorname{Tr}\bigg{[}\Phi\big{(}\rho\big{)}\bigg{]}=\operatorname{Tr}\bigg{(}\rho\bigg{)}.$ (2)
For completeness, once completely-positive maps and trace-preserving maps have been defined, I formally define CPTP maps.
Definition 5.
A completely positive trace-preserving map (i.e. a CPTP map) is a map that is both:
$\bullet$ Completely Positive.
$\bullet$ Trace-Preserving.
Modeling error as a CPTP map is justified as any map from and to density matrices is a CPTP map. Hence, any erroneous operation must be a CPTP map and so any erroneous implementation of an operation may be written as the errorless/ideal operation either followed or preceded by a CPTP map – as unitaries are CPTP maps.
Sec. II continues, from here, in Sec. II.1 with an introduction to the two characters of the problem setting: Alice and Bob. It explains their aims and capabilities, but that does not completely characterize the problem setting as their exact mode of interaction to achieve these aims are not yet specified. This requires an intermission, in Sec. II.2, from the presentation of the problem setting; Sec. II.2 defines the concepts of redaction and CPTP lists, which are vital for when the presentation of the problem setting is completed in Sec. II.3. Sec. II.3 specifies exactly how Alice and Bob interact to achieve their respective – and competing – aims, with the help of a new character, Robert, who is impartial and only wants to facilitate the interactions of Alice and Bob.
II.1 Introduction to Alice, Bob, and Robert
With its foundations established, I now present my problem setting – which is an adaption of the cryptographic setting [22] – beginning with the characters of the problem setting: Alice, Bob, and Robert.
In the adapted cryptographic setting used herein, Alice is attempting to get the result of a specific sampBQP [26, 27, 28] quantum computation [^5] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). of her choosing, while only having the ability to perform polynomially-bounded classical computation and initiate – and participate in – Protocol 1.
Bob is computationally unbounded and – using his participation in Protocol 1 when Alice invokes it – aims to trick Alice into accepting the results of an incorrect computation, believing they are the outputs of the computation she wanted to be performed, when they are not.
Bob represents the noise in the computation (the choices he makes in Protocol 1 are choosing which error get applied to the computation Alice requests) and is the adversary that “adversarial noise” gets its name from. This personification of the noise is used as a worst-case scenario, as real noise is not actually that smart or malicious, but if a protocol allows Alice to defend against Bob – in the adversarial problem setting – it will also work when computations experience the less sophisticated noise that is more typical in reality.
The limits of Bob are contained within Protocol 1 but these limits are enforced and personified within the modified adversarial problem setting by Robert, who also plays a role in Protocol 1 and is the one who actually performs the computations – that Alice is provided the results of – in the problem settings. Robert also checks that Alice and Bob are conforming to all the rules of Protocol 1, and aborts the protocol if not.
II.2 Problem Setting Preliminaries and Definitions
While the aims of Alice and Bob remain exactly the same as in the standard cryptographic setting, my specific problem setting is slightly modified from the traditional cryptographic setting. Mainly through limitations on Bob and the associated addition of a new, impartial character called Robert.
These limitations are entirely contained in how the circuits Alice requests be executed – and Bob tries to corrupt – are executed, which is presented in Protocol 1 in Sec. II.3. Protocol 1 features computations being performed not by Alice or Bob but by some third, honest, referee (whom I call Robert, as in Ref. [29]); with inputs (as prescribed by Protocol 1) from Alice and Bob. Alice providing the computations and Bob contributing the error. The choices of these inputs, given to Robert, are how Alice and Bob each attempt to achieve their respective goals. I.e. Alice requests computations; to Robert, in Protocol 1; and Bob chooses the errors to add to the computations during their execution, which Robert dutifully applies (assuming they conform to the requirements of Protocol 1) to the computations Alice requested.
Presenting Protocol 1 first requires I define a number of concepts. These start with Def. 6, Def. 7, and Def. 8. Which, collectively, define a concept called redaction, and its usage, which is designed to model – in the problem setting – the idea that noise is independent of which single-qubit gates are applied in a circuit but can depend on all other aspects of a circuit. As noise is adversarial in this paper, to model this I need some way of showing the adversarial noise (i.e. Bob) the circuit but hiding the single-qubit gates from it/him. Redaction is how I achieve this. It is not a physical thing we actually can do but is part of the new cryptographic model I develop herein – and represents single-qubit gates experiencing gate-independent error (as the gates are hidden when the CPTP maps applying the error are chosen by Bob).
II.2.1 Redaction and Related Definitions
I now commence the series of definitions required to adequately define redaction.
Definition 6.
A gate in a circuit is said to be redacted if the gate’s position in the circuit (i.e. when it is applied and to which qubits) is specified but the operator it represents (e.g. if it is a Pauli $\mathcal{X}$ gate or Pauli $\mathcal{Z}$ gate) is not. For an example of a circuit with a redacted gate, see Fig. 1.
Figure 1: An example circuit with a redacted gate (the gate with a black block, in the middle of the circuit). Note that the location of the gate (in terms of when it is applied and which qubits are affected by it) is depicted but which operation the gate represents is hidden by the black block and hence is unknown to anyone viewing this redacted circuit.
Similarly, a circuit is redacted (i.e is a redacted circuit) if all of its single-qubit gates are redacted, e.g. the circuit in Fig. 1.
Definition 7.
@@ II.2.1 Redaction and Related Definitions
Given a redacted circuit, the set of all circuits, without redactions, that the given redacted circuit possibly could be, if its redactions were removed, is called its redaction class. For example, all the circuits in Fig. 2 are in the redaction class of redacted circuit in Fig. 1.
Definition 8.
Two circuits are in the same redaction class if there exists a redacted circuit such that both the given circuits are in that redacted circuit’s redaction class.
For example, all the circuits in Fig. 2 are in the same redaction class: the redaction class of the redacted circuit in Fig. 1.
@@
Figure 2: Example circuits within the same redaction class, where the redaction class corresponds to the redacted circuit in Fig. 1. Note that the circuit in Fig. 1 could be any of the above circuits if the redaction on its single-qubit gate were removed.
II.2.2 CPTP Lists and Error-Related Definitions
@@
The second and final round of terminology required for the presentation of Protocol 1 is related to how Bob specifies the CPTP maps that implement the error in a specific execution of a circuit.
When Bob wants to provide Robert with CPTP maps that implement the error in a specific execution, he gives Robert a CPTP list, as defined in Def. 9. But Bob cannot just give Robert any CPTP list he likes, there are limits imposed by Protocol 1 (and rigidly enforced by Robert) that the CPTP lists Bob provides must conform to:
  1. 1.
    The CPTP list must match (as defined in Def. 10) the circuit it will be applied in the execution of (basically meaning the CPTP map is capable of defining the error in the circuit execution).
  2. 2.
    All CPTP lists Bob provided in a single instance of Protocol 1 must be from a single Set of Probabilistically Similar CPTP Lists with parameter $\beta$ (defined in Def. 12), declared at the beginning of Protocol 1, for a $\beta\in\mathbb{R}$ known before the start of the protocol.
@@
However, a CPTP list is just a list of CPTP maps. In order to to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps and each must act on exactly the right number of qubits. In this case the CPTP list is said to fit the circuit, as in Def. 10.
@@
A CPTP list fits a circuit if it may be used to describe the error [^6] Any resulting error is valid (from the perspective of fitting the circuit). It need not be in any way physically justified or correspond to a particular quantum computer’s typical error. in an execution of that circuit.
A CPTP list achieves this by containing exactly one CPTP map corresponding to each location in the circuit where error may occur [^7] These locations are: immediately after each gate, immediately after state preparation, and immediately before measurement. and each CPTP map acts on the required number of qubits (determined by the location it corresponds to in the circuit).
@@
For an example of a CPTP list fitting a circuit and then determining the error in a circuit execution, see Fig. 3.
@@
Note that notation is slightly abused to display the CPTP maps from $\Xi$ as gates (depicting error) in a circuit. To remedy this abuse slightly CPTP maps are highlighted in red to denote they are not unitaries but are CPTP maps.
@@
An important feature about CPTP lists and the circuits they fit is mentioned in Lemma 1. Proof is omitted as Lemma 1 follows from the above discussion.
@@
The error in a circuit execution is entirely determined by the CPTP list used to describe its error [^8] And – potentially – the outcomes of any stochastic processes in the application of the error it describes..
A key metric of error occurring in a circuit execution is defined in Def. 11, which allows for a useful quantification of the effect of the error.
@@
For any execution of a circuit afflicted by only stochastic Pauli error, the probability of error of that execution is the probability that the execution does not provide a sample from the same probability distribution that executing the circuit as intended (i.e. with no error applied) would.
I.e. it is the probability that none of the stochastic Pauli error channels in the circuit execution apply a Pauli gate. For example, if, for all density matrices, $\rho$:
$\xi_{1}(\rho)$ $=\dfrac{1}{2}\rho+\dfrac{1}{2}\hat{X}\rho\hat{X}^{\dagger}$ (3)
$\xi_{2}(\rho)$ $=\dfrac{1-0.01}{2}\rho+\dfrac{1+0.01}{2}\hat{X}\rho\hat{X}^{\dagger}$ (4)
$\xi_{3}(\rho)$ $=\dfrac{1+0.01}{2}\rho+\dfrac{1-0.01}{2}\hat{X}\rho\hat{X}^{\dagger},$ (5)
[…]
Having defined the probability of error in an execution of a circuit, which is a consequence of the CPTP list determining the error in that execution, I can define a formal concept, in Def. 12, wherein a set of CPTP maps that may govern error in circuit executions all have a similar probability of error.
@@
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (abbreviated as a SPSCLβ) is a set of CPTP lists, that all fit the same set of circuits (as in Def. 10), such that, for any CPTP list in the set, the probability of error in any circuit execution it determines the error for, once it is twirled to stochastic Pauli error, is within the interval:
@@
for some $P_{0}\in[0,1]$ (such that $P_{0}\big{(}1+\beta\big{)}\leq 4/5$).
For an example of a SPSCLβ, consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 11). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as an SPSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this SPSCLβ is $0.5$).
II.2.3 Relating CPTP Lists and Redaction Classes
With both SPSCLβ and redaction classes defined, they may be related to each other via the below Lemma 2. As will be seen, for each use of Protocol 1, Alice chooses a redaction class (implicitly, as all circuits must come from the same redaction class), and Bob chooses an SPSCLβ to determine the error in those executions. Hence, SPSCLβ and redaction classes are in some sense duals of each other. This duality is further expounded apon by Lemma 2.